# Survey on Modeling of Articulated Objects

### v1

@@ Abstract
3D modeling of articulated objects is a research problem within computer vision, graphics, and robotics. Its objective is to understand the shape and motion of the articulated components, represent the geometry and mobility of object parts, and create realistic models that reflect articulated objects in the real world. This survey provides a comprehensive overview of the current state-of-the-art in 3D modeling of articulated objects, with a specific focus on the task of articulated part perception and articulated object creation (reconstruction and generation). We systematically review and discuss the relevant literature from two perspectives: geometry processing and articulation modeling. Through this survey, we highlight the substantial progress made in these areas, outline the ongoing challenges, and identify gaps for future research. Our survey aims to serve as a foundational reference for researchers and practitioners in computer vision and graphics, offering insights into the complexities of articulated object modeling.
volume: 43issue: 2
@@ 1 Introduction
Articulated objects, composed of multiple rigid parts connected by joints, are ubiquitous in our daily life, encompassing a wide range of objects, such as human bodies, mechanical assemblies, furniture pieces, etc. Modeling these objects is a research field intersecting computer vision, graphics, and robotics. It aims to comprehend and represent the shape and mobility of articulated components. This endeavor extends to creating 3D models that realistically reflect real-world articulated objects. Contributing to the dynamics and functionality in our physical world, articulated object modeling is essential for a wide range of applications, such as animation [YVN*22, CJS*23, QWM*23, LZWL23] and simulation [WCK19, YZW*22, XQM*20], robotic interaction and manipulation [HNOS15, GES21, MGM*21, QF23], embodied AI [KMH*17, PRB*18, GGS*19, SKM*19, PUS*23], etc.
The complexity of articulated object modeling stems from the fact that articulated objects are not just represented by the geometry of their parts but also by their kinematic structure. Over the past decade, much research has focused on addressing these complexities in two main directions: articulated part perception, and articulated object reconstruction and generation. Articulated part perception focuses on the detection and segmentation of the articulated components of an object while estimating the articulation model that describes the kinematic structure of the object. Articulated object creation involves the reconstruction and synthesis of these parts into a compositional and hierarchical 3D model that reflects the real-world object that can be interacted with.
In this survey, we begin by laying the groundwork with an overview of the field, defining the scope and focus of our discussion in Section 2. Next, in Section 3, we compile the key datasets for articulated objects that have been collected and utilized within the research community, highlighting their critical role in advancing the 3D modeling of articulated objects. Our survey then delves into an in-depth analysis of the techniques and methodologies developed for 3D modeling of articulated objects. This analysis is organized around two pivotal axes: geometry processing and articulation modeling, which are the foundational components to tackle the complexities in this field. Geometry processing covers the representation and approaches employed to understand the shape of the articulated parts, whereas articulation modeling involves the articulation model and the methodologies used to estimate the part mobility and kinematic structure of the object. In the end, we conclude our survey by identifying the key challenges and potential research directions that can drive future progress in the field.
@@
This survey provides a comprehensive overview of the current state-of-the-art in 3D modeling of articulated objects. Through this survey, we highlight the substantial progress made in these areas, outline the ongoing challenges, and identify the gaps for future research. Our survey aims to serve as a foundational reference for researchers and practitioners in computer vision and graphics, offering insights into the complexities of articulated object modeling and inspiring new research in this area.
@@
Related surveys. The most related survey to ours is the recent survey by [PŠDC22, ], which focuses on the manipulation of articulated objects in the realm of robotics and computer vision. In contrast, our focus is on the geometry and articulation modeling of articulated objects within the field of computer vision and graphics. This aspect represents a distinct yet critical dimension of articulated object research, orthogonal to the topic of manipulation. Additionally, another related survey by [HSvK18, ] explores the functionality of general objects within the scope of computer graphics literature. The functionality of an object is influenced not just by its mobility but also by its shape, material, and various other attributes that collectively determine its capability to perform a task. Our survey takes a more specialized focus on the articulated object and its mobility particularly, compared to the work by [HSvK18, ] which concentrates on functionality for general objects.
@@
Figure 1: Common joint types. The first row shows the joints with linear motion which is the most common assumption considered in the literature. The second row shows more complex joints with more than 1 DoF. Figure reproduced from original paper [Mue19].
@@ 2.1 Definition of Articulated Objects
An articulated object is composed of multiple rigid parts interconnected by joints. These joints serve as the pivotal points that allow for relative motion between the connected parts. The range and nature of the articulation of each part are constrained and defined by the type of joint it possesses. Among man-made objects, the most frequently encountered joint types facilitate linear motion, such as revolute, prismatic, and helical joints. For example, a laptop is an articulated object with a revolute joint connecting the screen to the keyboard. This joint allows the screen to rotate around a single axis relative to the keyboard, with the range of motion being restricted to a specific limit. In contrast, the human body is a more complex example of an articulated object, with a wider variety of joint types, such as ball-and-socket or spherical joints. These joints allow for motion beyond linear translation or rotation, enabling a more diverse range of movements. The illustration of several common joint types is shown in Figure 1.
@@
  • Organic objects refers to naturally occurring systems that are capable of initiating movement through articulated structures. The bodies of humans and other animals are common examples of organic articulated objects. They are crucial for various functions such as locomotion, manipulation, and interaction with the environment. A skeleton with bones and joints is usually used to represent the kinematic structure of these objects. The motion allowed for these objects is usually structurally complex with a relatively large number of joints varied in the topology of connections and degrees of freedom. However, the kinematic structure is relatively fixed and can be predefined for a specific species. How to extract effective skeletons, reconstruct 3D models, and fit dynamic motions for animals across various species is an active research topic [WLJ*23, AM23, LWP*23, LLL*24, YRH*24]. Leveraging the skeleton template of the human body, and coordination between the joints to achieve the desired animation with high fidelity is another line of research [SYZR21, BKY*22, SBR22, WSGT22, LZWL23, QWM*23]. In this survey, we leave the discussion of organic articulated objects out of the scope and only focus on human-made articulated objects.
@@
  • Human-made objects are engineered systems designed to mimic the natural articulation of organic objects or to serve specific functionalities that require movement. These objects are complex assemblies of rigid parts connected by joints, and their motion is often a direct result of interaction with the environment or driven by an active system. Common examples of human-made articulated objects include gadgets, furniture, vehicles, and other mechanical systems. These objects are ubiquitous in our daily lives and play an important role in our interaction with the physical world. Modeling human-made articulated objects presents several unique challenges in terms of physical and geometric modeling, articulation modeling, and environmental interaction. The modeling of these objects is the focus of this survey and will be discussed in detail in the following sections. We will only mention articulated objects for short in the rest of the survey to refer to human-made articulated objects.
Figure 2: Common geometric representation for articulated object modeling. Figures reproduced from original papers [HLV*17, YHY*19].
@@ 2.2 Representation of Articulated Objects
Articulated object modeling is a multifaceted domain within computer vision, graphics, and robotics. It aims to understand the shape and interactability of the articulated components, represent the geometry and mobility of these objects, and create realistic models that reflect articulated objects in the real world. The inherent complexity of this task arises from the fact that the articulation of the object is not only determined by the geometry of the parts but also by the kinematic structure of the object. As a result, effective modeling of articulated objects requires the capturing of a dual representation of the geometry and articulation of the object. These two facets are deeply interconnected, each influencing and informing the other. The geometric decomposition of the parts lays the foundational groundwork for analyzing part mobility, while the process of articulation modeling also guides the shape understanding of the mobility parts. This symbiotic relationship between geometric and articulation modeling underscores the importance of a cohesive approach in articulated object modeling, where understanding and representing both geometry and articulation are essential for capturing the full essence and functionality of complex entities. Over the last decade, there has been a notable surge in research within this field, addressing the inherent challenges predominantly from two perspectives: geometry processing and articulation modeling. We will further discuss these two axes in the following sections.
@@
Geometric representation. Unlike the geometric modeling of static objects, which typically concentrates solely on the object’s outer surface, the geometric representation of articulated objects demands a more nuanced approach. This involves capturing the spatial arrangement of the parts, modeling the part-level surfaces, and constructing the interior structure of the object. It is essential to accurately describe the shape, size, relative position, and orientation of each individual part that is structurally organized within the object. This comprehensive description is vital as it provides the foundational details needed to represent a complete object that can be interacted with. The data format chosen for representing the geometry of the objects varies, depending on data capture methods and the intended application. The most common access to the part geometry is through the 3D point cloud, which is a set of points sampled from the object surface that can be obtained from 3D scanning or depth projection. While the point cloud provides a raw, unstructured representation of the object’s geometry, it often undergoes further processing into more organized formats for detailed analysis. Another commonly employed representation is the 3D mesh, which consists of a collection of polygons that approximate the surface of each part. Due to its structured nature and ability to effectively model surface details, the mesh format is extensively utilized in applications involving simulation, rendering, and animation. In Section 4, we will discuss different choices of geometric representation in various related tasks in detail. Here we specify the geometric representations that are considered in the literature and visually illustrated in Figure 2.
@@
  • Mesh: is a collection of polygons that approximate the surface.
@@
  • Implicit field: is an implicit function that assigns a value to each point in the space to construct a field describing the object shape. Common forms of implicit fields include the signed distance field (SDF), occupancy field, density field, etc. The function parameterized by a neural network, known as a neural field, is a popular intermediate representation as it is inherently differentiable and easy to integrate with learning-based methods.
  • 2D images: some work approaches the 3D modeling problem from the 2D image domain, such as single-view RGB-D images and posed RGB images captured from multi-views or stereo cameras.
Figure 3: Common articulated motion representation. Figures reproduced from original papers [YHL*18, JLCN21, SHL*14].
Articulation representation. The representation for articulation includes the description of the mobility of each part and the kinematic structure of the object. The kinematic structure describes how the object’s parts are connected and how they can move relative to each other. Typically, this kinematic structure is represented by a tree or graph, which is an effective way to model hierarchical relationships in articulated systems. The part mobility can described by the parameters of each connected joint or the deformation field of the object. In Section 5, we will discuss ways of articulation modeling with different assumptions and representation in various related tasks in detail. Here we specify the articulated motion representation as follows and visually illustrated in Figure 3.
@@
  • Joint parameters: the parameters of each joint, including the joint type, location and orientation of the joint axis, the rotational angle or translational distance as joint state, and the left/right bound of joint as joint limit or motion range.
@@
[HLV*17, ] synthetic object mesh 368 368 -
RPM-Net [YHY*19] synthetic object mesh 969 1,420 -
Shape2Motion [WZS*19] synthetic object mesh 2,440 6,762 45
RBO [MEB19] synthetic, real object mesh, RGB-D 14 21 14
PartNet-Mobility [XQM*20] synthetic object mesh 2,346 14,068 46
ReArt-48 [LXX*22] real object mesh 48 ? 5
AKB-48 [LXF*22] real object mesh 2,037 ? 48
OPDSynth [JMSC22] synthetic object mesh, RGB 683 1,343 11
OPDReal [JMSC22] real object mesh, RGB-D 284 875 8
MultiScan [MZJ*22] real scene mesh, RGB-D 10,957 5,129 20
GAPartNet [GXZ*23] synthetic, real object mesh 8,489 1,166 27
Table 1: Datasets for articulated objects with statistics on the source of collection, the level at which these datasets operate, data representation, whether the 3D model is texture, the number of objects, movable parts, and object categories. The ’-’ symbol indicates that the information is not available, and the ’?’ symbol indicates that the information is not reported in the original paper.

### v3

@@ Abstract
3D modeling of articulated objects is a research problem within computer vision, graphics, and robotics. Its objective is to understand the shape and motion of the articulated components, represent the geometry and mobility of object parts, and create realistic models that reflect articulated objects in the real world. This survey provides a comprehensive overview of the current state-of-the-art in 3D modeling of articulated objects, with a specific focus on the task of articulated part perception and articulated object creation (reconstruction and generation). We systematically review and discuss the relevant literature from two perspectives: geometry modeling (i.e., structure and shape of articulated parts) and articulation modeling (i.e., dynamics and motion of parts). Through this survey, we highlight the substantial progress made in these areas, outline the ongoing challenges, and identify gaps for future research. Our survey aims to serve as a foundational reference for researchers and practitioners in computer vision and graphics, offering insights into the complexities of articulated object modeling.
volume: 44issue: 2
@@ 1 Introduction
Articulated objects, composed of multiple rigid parts connected by joints, are ubiquitous in our daily life, encompassing a wide range of objects, such as mechanical assemblies, furniture pieces, even human bodies and other animals. Modeling these objects is a research field intersecting computer vision, graphics, and robotics. It aims to comprehend and represent the shape and mobility of articulated components. This endeavor extends to creating 3D models that realistically reflect real-world articulated objects. Contributing to the dynamics and functionality in our physical world, articulated object modeling is essential for a wide range of applications, such as animation [yang2022banmo, chen2023fast, qian20233dgs, li2023animatable] and simulation [weng2019photo, yang2022object, xiang2020sapien], robotic interaction and manipulation [hausman2015active, gadre2021act, mo2021where2act, qian2023understanding], embodied AI [kolve2017ai2, puig2018virtualhome, gao2019vrkitchen, savva2019habitat, puig2023habitat], etc.
The complexity of articulated object modeling stems from the fact that articulated objects are not just represented by the geometry of their parts but also by their kinematic structure. Over the past decade, much research has focused on addressing these complexities in two main directions: articulated part perception, and articulated object creation. The primary goal of articulated part perception is to analyze the mobility of parts in an object, which is useful in two application areas: 1) object manipulation in robotics; and 2) interaction and animation in simulations. The vision and robotics communities have been actively working on the first application area [gadre2021act, mo2021where2act, eisner2022flowbot3d, xu2022universal, wang2022adaafford, mittal2022articulated, geng2023partmanip, bao2023dexart, hsu2023ditto, ning2024where2explore, ling2024articulated, yu2024gamma, wang2024rpmart], where the task is usually decomposed to affordance prediction, articulation estimation, and action planning. We leave this line of work out of the scope of this survey, as it is focused on manipulation of articulated objects with robots in the real world. Another line of work from vision and graphics community focuses more on the geometric and kinematic modeling of the articulated objects for the second application area, where the task is usually decomposed to articulated part segmentation and kinematic structure understanding. For articulated object creation, the main goal is to reconstruct or generate articulated parts into a compositional and hierarchical 3D model that represents a real-world object that can be interacted with in a virtual environment.
@@
This survey provides a comprehensive overview of the current state-of-the-art in 3D modeling of human-made articulated objects. Through this survey, we highlight the substantial progress made in these areas, outline the ongoing challenges, and identify the gaps for future research. Our survey aims to serve as a foundational reference for researchers and practitioners in computer vision and graphics, offering insights into the complexities of articulated object modeling and inspiring new research in this area.
We begin by laying the groundwork with an overview of the field, defining the scope and focus of our discussion in Section 2. Next, in Section 3, we compile the key datasets for articulated objects that have been collected and utilized within the research community, highlighting their critical role in advancing the 3D modeling of articulated objects. Our survey then delves into an in-depth analysis of the techniques and methodologies developed for 3D modeling of articulated objects. This analysis is organized around two pivotal axes: geometry modeling and articulation modeling, which are the foundational components to tackle the complexities in this field. Geometry modeling covers the representation and approaches employed to understand the shape of the articulated parts, whereas articulation modeling involves the articulation model and the methodologies used to estimate the part mobility and kinematic structure of the object. In the end, we conclude our survey by identifying the key challenges and potential research directions that can drive future progress in the field.
@@
Related surveys. The most related survey to ours is the recent survey by pejic2022articulated, which focuses on the manipulation of articulated objects in the realm of robotics and computer vision. In contrast, our focus is on the geometry and articulation modeling of articulated objects within the field of computer vision and graphics. This aspect represents a distinct yet critical dimension of articulated object research, orthogonal to the topic of manipulation. Additionally, another related survey by hu2018functionality explores the functionality of general objects within the scope of computer graphics literature. The functionality of an object is influenced not just by its mobility but also by its shape, material, and various other attributes that collectively determine its capability to perform a task. Our survey takes a more specialized focus on the articulated object and its mobility particularly, compared to the work by hu2018functionality which concentrates on functionality for general objects.
@@
Figure 1: Common joint types. The first row shows the joints with linear motion which is the most common assumption considered in the literature. The second row shows more complex joints with more than one DoF. Figure reproduced from original paper lynchpark2017modern.
@@ 2.1 Definition of Articulated Objects
An articulated object is composed of multiple rigid parts interconnected by joints. These joints serve as the pivotal points that allow for relative motion between the connected parts. The range and nature of the articulation of each part are constrained and defined by the type of joint it possesses. Among human-made objects, the most frequently encountered joint types facilitate linear motion, such as revolute, prismatic, and helical joints. For example, a laptop is an articulated object with a revolute joint connecting the screen to the keyboard. This joint allows the screen to rotate around a single axis relative to the keyboard, with the range of motion being restricted to a specific limit. In contrast, the human body is a more complex example of an articulated object, with a wider variety of joint types, such as ball-and-socket or spherical joints. These joints allow for motion beyond linear translation or rotation, enabling a more diverse range of movements. The illustration of several common joint types is shown in Figure 1.
@@
  • Organic objects refer to naturally occurring systems that are capable of initiating movement through articulated structures. The bodies of humans and other animals are common examples of organic articulated objects. They are crucial for various functions such as locomotion, manipulation, and interaction with the environment. A skeleton with bones and joints is usually used to represent the kinematic structure of these objects. The motion allowed for these objects is usually structurally complex with a relatively large number of joints varied in the topology of connections and degrees of freedom. However, the kinematic structure is relatively fixed and can be predefined for a specific species. How to extract effective skeletons, reconstruct 3D models, and fit dynamic motions for animals across various species is an active research topic kulkarni2020articulation; yang2021lasr; wu2022casa; yao2022lassie; wu2023magicpony; aygun2023saor; lei2023gart; yao2023hi; li2024learning; yao2024artic3d; liu2024lepard; jakab2024farm3d; maheshwari2023transfer4d; yang2024omnimotiongpt. Leveraging the skeleton template of the human body or hands and coordination between the joints to achieve the desired animation aberman2020skeleton; su2021nerf; bergman2022generative; su2022danbo; wang2022arah; kuai2023camm; lei2023gart; zhang2023generating; li2023animatable; qian20233dgs; tan2023distilling or human-object interaction taheri2020grab; zhang2020perceiving; xu2021d3d; cao2021reconstructing; zhang2022couch; fan2022articulated; bhatnagar2022behave; qian2022understanding; ghosh2023imos; jiang2023full; huang2023one; lee2023locomotion; xie2024rhobin is another line of research. We direct readers to a recent survey on human body modeling berretti2018representation; chen2021towards; wang2021deep; mourot2022survey; akber2023deep; zheng2023deep; yang2024digital and 3D hand modeling cheng2015survey; wheatland2015state; barsoum2016articulated; li2019survey; ahmad2019hand for more details. In this survey, we leave the discussion of organic articulated objects out of the scope and only focus on human-made articulated objects.
@@
  • Human-made objects are engineered and designed to mimic the natural articulation of organic objects or to serve specific functionalities that require movement. These objects are complex assemblies of rigid parts connected by joints, and their motion is often a direct result of interaction with the environment or driven by an active system. Common examples of human-made articulated objects include gadgets, furniture, vehicles, and other mechanical systems. These objects are ubiquitous in our daily lives and play an important role in our interaction with the physical world. Modeling human-made articulated objects presents several unique challenges in terms of physical and geometric modeling, articulation modeling, and environmental interaction. Modeling these objects is the focus of this survey and will be discussed in detail in the following sections. We will only mention articulated objects for short in the rest of the survey to refer to human-made articulated objects.
Key differences. Modeling organic and human-made articulated objects presents distinct challenges due to differences in geometry, kinematic structure, motion patterns, and interaction with the environment. Topologically, organic objects share consistent skeletal structures within species, enabling predefined templates (e.g., human body, quadrupeds), whereas human-made objects exhibit greater variability. Some categories follow standard designs (e.g., eyeglasses, scissors), while others (e.g., storage furniture) vary significantly across instances. In terms of motion, organic objects have biomechanical joints driven by muscle forces, leading to soft-tissue deformations, whereas human-made objects typically follow rigid-body mechanics with constrained degrees of freedom. While this generally simplifies articulation modeling, challenges arise in accurately estimating joint parameters in relation to object geometry, handling nested joints with complex kinematic dependencies (e.g., bi-fold doors, folding umbrellas), and accounting for real-world imperfections such as frictional resistance and mechanical backlash. Additionally, organic objects evolve for locomotion and behaviors, while human-made objects are designed for specific functionalities, influencing how they articulate and are manipulated. Consequently, research on organic objects emphasizes geometric reconstruction and animation, whereas human-made objects are studied in the context of 3D modeling for robotic manipulation and functional interaction.
These fundamental differences make it difficult for methods developed for one type to directly generalize to the other. Structure-wise, the parametric models and statistical priors used for organic objects rely on consistent skeletal structures, which are not scalable to accommodate the diverse topologies of human-made objects. Geometrically, organic objects are typically represented as a single continuous surface, i.e., skin enclosing a skeleton. Reconstruction methods optimized for organic shapes rely on this assumption and are not well-suited for human-made objects, which often consist of multiple discrete, interlocking parts with internal structures, requiring different modeling approaches. Articulation-wise, organic objects exhibit complex, non-linear motion patterns driven by muscle forces, necessitating biomechanical models and motion capture data for accurate representation. On the other hand, the rigid-body mechanics of human-made objects are more amenable to analytical kinematic models, simplifying articulation modeling but requiring precise joint parameter estimation to enable plausible interaction in physical simulations.
@@ 2.2 Representation of Articulated Objects
Articulated object modeling is a multifaceted domain within computer vision, graphics, and robotics. It aims to understand the shape and interactability of the articulated components, represent the geometry and mobility of these objects, and create realistic models that reflect articulated objects in the real world. The inherent complexity of this task arises from the fact that the articulation of the object is not only determined by the geometry of the parts but also by the kinematic structure of the object. As a result, effective modeling of articulated objects requires capturing a dual representation of the geometry and articulation of the object. These two facets are deeply interconnected, each influencing and informing the other. The geometric decomposition of the parts lays the foundational groundwork for analyzing part mobility, while the process of articulation modeling also guides the shape understanding of the mobility parts. This symbiotic relationship between geometric and articulation modeling underscores the importance of a cohesive approach in articulated object modeling, where understanding and representing both geometry and articulation are essential for capturing the full essence and functionality of complex entities. Over the last decade, there has been a notable surge in research within this field, addressing the inherent challenges predominantly from two perspectives: geometry modeling and articulation modeling. We will further discuss these two axes in the following sections.
Figure 2: Common geometric representations for articulated object modeling. Figures reproduced from original papers liu2024cage; yan2019rpm; schmidt2014dart.
@@
Geometric representation. Unlike the geometric modeling of static objects, which typically concentrates solely on the object’s outer surface, the geometric representation of articulated objects demands a more nuanced approach. This involves capturing the spatial arrangement of the parts, modeling the part-level surfaces, and constructing the interior structure of the object. It is essential to accurately describe the shape, size, relative position, and orientation of each individual part that is structurally organized within the object. This comprehensive description is crucial, as it provides the foundational details necessary to represent a complete object that can be interactively manipulated. The data format used to represent object geometry varies based on the data capture methods and the intended application. The most common access to the part geometry is through the 3D point cloud, which is a set of points sampled from the object surface that can be obtained by 3D scanning or depth projection. While point clouds provide a raw, unstructured representation of the object’s geometry, they often undergo further processing into more organized formats for detailed analysis. Another common representation is 3D mesh, which approximates the surface of each part with primitives such as polygons. The mesh format is widely used in applications such as simulation, rendering, and animation due to its ability to effectively model surface details and because it can flexibly represent object structure. In Section 4, we will discuss different choices of geometric representation in various related tasks in detail. Here we specify the geometric representations that are considered in the literature and visually illustrated in Figure 2.
@@
  • Mesh: is a collection of vertices, edges, and polygons usually with a structured connectivity (e.g., 2-manifold) that approximate the surface.
@@
  • Implicit field: is an implicit function that assigns a value to each point in the space to construct a field describing the object. Common forms of implicit fields include the signed distance field (SDF), occupancy field, density field, etc. The function parameterized by a neural network, known as a neural field, is a popular intermediate representation as it is inherently differentiable and easy to integrate with learning-based methods.
  • Images-based: some work approaches the 3D modeling and perception problem from the image-based domain, such as depth images, RGB-D images and posed RGB images captured from multi-views or stereo cameras.
Figure 3: Common articulated motion representations. Figures reproduced from original papers yi2018deep; liu2024singapo.
Articulation representation. The representation for articulation includes the description of the mobility of each part and the kinematic structure of the object. The kinematic structure describes how the object’s parts are connected and how they can move relative to each other. Typically, this kinematic structure is represented by a tree or graph, which is an effective way to model hierarchical relationships in articulated systems. The part mobility can be described by the parameters of each connected joint or the deformation field of the object. In Section 5, we will discuss ways of articulation modeling with different assumptions and representation in various related tasks in detail. Here we specify the articulated motion representation as follows and visually illustrated in Figure 3.
@@
  • Joint parameters: the parameters of each joint, including the joint type, the joint axis, the joint state such as the rotational angle or translational distance, and the motion range or joint limit.
@@
Figure 4: Representative datasets for 3D articulated objects: PartNet-Mobility xiang2020sapien for synthetic data, AKB-48 liu2022akb for real scans of objects, MultiScan mao2022multiscan for real indoor scenes with articulated objects, and GAPartNet geng2023gapartnet for synthetic and real objects with fine-grained annotations for the generalizable actionable parts. ParaHome kim2024parahome is a recent dataset providing human-object interactions with articulated objects.
@@
hu2017learning synthetic object mesh 368 368 -
RPM-Net yan2019rpm synthetic object mesh 969 1,420 43
Shape2Motion wang2019shape2motion synthetic object mesh 2,440 6,762 45
RBO martin2019rbo synthetic, real object mesh, RGB-D 14 21 14
PartNet-Mobility xiang2020sapien synthetic object mesh 2,346 14,068 46
ReArt-48 liu2022toward real object mesh 48 - 5
AKB-48 liu2022akb real object mesh 2,037 - 48
OPDSynth jiang2022opd synthetic object mesh, RGB 683 1,343 11
OPDReal jiang2022opd real object mesh, RGB-D 284 875 8
MultiScan mao2022multiscan real scene mesh, RGB-D 10,957 5,129 20
OPDMulti sun2023opdmulti real scene RGB-D 217 688 33
GAPartNet geng2023gapartnet synthetic, real object mesh 8,489 1,166 27
ACD iliash2024s2o synthetic object mesh 354 1,350 21
ParaHome kim2024parahome real object mesh 22 - 8
Table 1: Datasets related to articulated objects with statistics on the source of collection, the level at which these datasets operate, data representation, whether the 3D model is texture, the number of objects, movable parts, and object categories. The ’-’ symbol indicates that the information is not available or not reported in the original paper.